how do you solve this system by substitution then how the solution would be graph? thanks.
y=x^3+x^2+x
y=-3x

Answer by edjones(7569)

y=x^3+x^2+x
y=-3x
.
x^3+x^2+x=-3x
x^3+x^2+4x=0
x(x^2+x+4)=0
x=0
.
x^2+x+4=0
x^2+x =-4
x^2+x+ 1/4=-4+ 1/4 Complting the square.
(x + 1/2)^2= -15/4
x + 1/2 = +-sqrt(-15)/2
x=-1/2 +- sqrt(15)i /2
x=(-1+sqrt(15)i)/2, x=(-1-sqrt(15)i)/2, x=0
.
Ed